Q if sec 4 Ѳ+ sec 2 Ѳ=10+ tan 4 Ѳ+tan 2 Ѳ then sin 2 Ѳ= ? the answer is 4/5 (4 divided b 5)

  1. Q   if sec4Ѳ+ sec2 Ѳ=10+ tan4 Ѳ+tan2 Ѳ then sin2 Ѳ= ?
  2. the answer is 4/5 (4 divided b 5)

Grade:11

1 Answers

Sunil Raikwar
9 years ago
sec4Ѳ+ sec2Ѳ=10+ tan4Ѳ+tan2Ѳ
sec4Ѳ – tan4Ѳ = 10+tan2Ѳ – sec2Ѳ
(sec2Ѳ+tan2Ѳ)(sec2Ѳ-tan2Ѳ) =10-1
sec2Ѳ+tan2Ѳ=9
1+tan2Ѳ+tan2Ѳ=9
2tan2Ѳ = 8
tan2Ѳ = 4
therefore sin2Ѳ=4/5